OpenStudy (anonymous):

Assume that the mean wait time spent by women in a restroom line at a sporting event is 6.3 minutes with a standard deviation of 1.5 minutes. For this distribution, find a raw score that corresponds to a z-score of 1.8

jimthompson5910 (jim_thompson5910):

z = (x - mu)/sigma 1.8 = (x - 6.3)/1.5 solve for x

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

x=8.930